the cartesian coordinates of a point are given by (x,y)=(-2.5,-3.5) meters. Find the polar coordinates of this point?



1
Expert's answer
2021-07-02T09:38:12-0400

The polar coordinates are related with cartesian coordinates by next equations

"\rho=\sqrt{x^2+y^2}, \quad \theta=\tan^{-1}\frac{y}{x}"

We get

"\rho=\sqrt{(-2.5)^2+(-3.5)^2}=4.3, \\
 \theta=\tan^{-1}\frac{-3.5}{-2.5}=234^{\circ}"